OpenStudy (anonymous):

an 18ft ladder leans against a wall. the bottom of the ladder is 5ft from the wall at t=0 and slides away at a rate of 2 ft/sec. Find the velocity at t=3

OpenStudy (anonymous):

unless I'm missing something, the velocity is 2ft/sec and it's constant, isn't it?

OpenStudy (anonymous):

i have to use derivatives it is a related rates problem

OpenStudy (anonymous):

Do you have any sample at all?

OpenStudy (anonymous):

what do you mean sample?

OpenStudy (anonymous):

z² = x² + y² ->y dy/dt = - x dx/dt => dy/dt = - x dx/dt / y At t = 3 => x = 5 ft + 3 * 2 = 11 ft => y = √ (18² - 11²) = 14.25 ft Thus dy/dt = - 11 * 2 / 14.25 = -1.54 ft/sec

OpenStudy (anonymous):

Your lecture or some homework?

OpenStudy (anonymous):

That's the best I can visualize! It'd be perfect if I know what are you learning!

OpenStudy (anonymous):

oh yea i had an example from lecture.. but i couldn't figure out from my notes where one of the numbers came from

OpenStudy (anonymous):

how can the answer be negative?

OpenStudy (anonymous):

Could you attach, so I can make sure my solution is 100% correct!

OpenStudy (anonymous):

im not sure how to attach but it is correct thank you

OpenStudy (anonymous):

Negative is the direction, which the laddter is sliding down!

OpenStudy (anonymous):

How do you know it's correct :)

OpenStudy (anonymous):

online homework

OpenStudy (anonymous):

So you enter the result already?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yeehh, great! I hate to give sloppy answer!

OpenStudy (anonymous):

The online homework is very tricky, sometimes if you're miss the sign or decimal point, your answer is rejected!